OpenStudy (anonymous):

Oxides of nitrogen are pollutant gases which are emitted from car exhausts. In urban traffic, when a car travels one kilometre, it releases 0.23 g of an oxide of nitrogen NxOy, which occupies 120 cm3. What are the values of x and y? (Assume 1 mol of gas molecules occupies 24.0 dm3.) A x = 1, y = 1 B x = 1, y = 2 C x = 2, y = 1 D x = 2, y = 4 the ans. is B.. can somebody explain the working?

OpenStudy (anonymous):

its easy you need to calculate molar mass of that compound NxOy by next steps: n=V/Vm = 0,12 dm3/24 dm3mol-1 = 0,005 moles n=m/M => M=m/n = 0,23g/0,005mol = 46 gmol-1 and now you see that only posibility is NO2 :D

OpenStudy (anonymous):

thnx